OpenStudy (anonymous):

Solve the equation: 1/3x - 4 = 20. A. x = 5.3 B. x = 8 C. x = 48 D. x = 72

OpenStudy (aperez):

So after you get the answer for from the 20 + 4 , do you like add that too 1/3x?

OpenStudy (anonymous):

umm i mean (x/3)-4=20??

OpenStudy (anonymous):

if 1/3x - 4 = 20, x=1/72 if x/3 - 4 = 20, x=72

OpenStudy (aperez):

Oh okay that's more understandable thanks :)